If X is the number scored in a throw of a fair die, show that the Chebychev’s inequality gives , P{|X-µ|2.5}<0.47, where µis the mean of X, while the actual probability is zero.
1
Expert's answer
2020-03-03T16:14:45-0500

When a die is thrown, "X" denotes the number that turns up. We find that the mean of "X" is


"\mu=E(X)={7 \over 2}, Var(X)=\sigma^2={35 \over 12}"

By Chebyshev's inequality we have


"P(|X-\mu|\geq C)\leq{\sigma^2 \over C^2}"

Taking "C=2.5," we have


"P(|X-\mu|>2.5)<{{35 \over 12} \over 2.5^2}"

"P(|X-\mu|>2.5)<0.47"

Or

"P(|X-3.5|>2.5)<0.47"